OpenStudy (anonymous):

please walk me through the steps to figure this out: y^2+x^2=53 and y-x=5

OpenStudy (superdavesuper):

u will end up with a quadratic eqn: ax^2 +bx + c = 0 do u know how to solve that?

OpenStudy (anonymous):

uhm no, thats just the quadratic formula right?

OpenStudy (superdavesuper):

yes...good that u know the quadratic formula....back to ur Q.... y-x=5 so by re-arranging, y=?

OpenStudy (anonymous):

y=5+x

OpenStudy (anonymous):

y^2(y-5)^2=53?

OpenStudy (superdavesuper):

Sure...think u missed a "+" somewhere though :) Expand (y-5)^2 and what will u get?

OpenStudy (anonymous):

y^2-25?

OpenStudy (anonymous):

oh yea i did! whoops, y^2+(y^2-25)=53

OpenStudy (superdavesuper):

Nope (y+a)^2 = y^2 + 2ay + a^2....so try that plz?

OpenStudy (anonymous):

what is a ?

OpenStudy (superdavesuper):

scroll up see what u did...a in this case is -5 right?

OpenStudy (anonymous):

(y-5)^2=y^2+2(-5)y+5^2 is that what you mean

OpenStudy (superdavesuper):

yup yup n put that back into ur original eqn plz?

OpenStudy (anonymous):

y^2+y^2+2(-5)y+5^2=53?

OpenStudy (superdavesuper):

rearrange n regroup....I will do one part for you: +5^2-53 is -28

OpenStudy (superdavesuper):

plz do the same for y^2 and y

OpenStudy (anonymous):

i got it differently. 25+10x+2x^2=53

OpenStudy (anonymous):

-28+10x+2x^2=0

OpenStudy (superdavesuper):

Hmmmm what happens to the y? :P so the eqn now looks like 2y^2 + 10y + (-28) = 0 right?